# Understanding Teen Overreliance on AI Companion Chatbots Through Self-Reported Reddit Narratives

### v3

@@ 3.1. Data Collection, Preprocessing, and Filtering
We scraped posts from the r/CharacterAI subreddit (with over 2.5 million active members  [^4] https://www.reddit.com/r/CharacterAI/) using the PullPush Reddit Search API [^5] https://pullpush.io/#docs from January 1, 2023, to April 8, 2025. To scope the data, we queried this dataset with a list of keywords that related to addiction and overreliance: addict*, dependant, dependent, obsess*, attach*, rely*. While we refrain from using the term addict, we recognized that teens used this term when describing their overreliance and included it in our keyword search. We refined our query by reviewing posts to determine if these keywords resulted in finding relevant posts. We expanded our search with additional identified keywords that recurred among these posts, such as quit, social life, craving, unhealthy, escaping, and one key phrase ”I can’t live without”. After collecting posts using keywords which resulted in 6,133 posts, we removed 746 posts, including 428 duplicates, and 170 marked as “removed” or “deleted.” After this process 5,535 posts remained.
Since our goal was to study this issue from the perspective of teens aged 13 to 17, we first needed a way to identify whether a post was written by a teenager. We could not use user profile information for this, since most Reddit accounts in our dataset lacked identifiable details such as age. We therefore relied only on the content of posts.
Large Language Models (LLMs) have shown strong performance in zero-shot settings, sometimes even outperforming traditional machine learning and deep learning methods in text classification tasks (Wang et al., 2023). Motivated by this, we used OpenAI’s GPT-4o mini model as a classifier to identify whether a post included relevant age disclosure for the 13–17 age range. The following text is a summarized version of the prompt used to guide the model (the full version appears in Appendix A):
@@
To evaluate the model’s consistency with human judgment, the first author annotated a subset of 100 random posts as either True (post made by a teen) or False (post not made by a teen). We calculated the Inter-Rater Reliability (IRR) (Cohen’s kappa = 0.74) between the author’s and the model’s annotations. We chose IRR instead of conventional classification performance metrics, such as accuracy, as these metrics are highly sensitive to class imbalance (Mullick et al., 2020). Since the true distribution of teen vs. non-teen posts in the dataset was unknown and likely imbalanced, and our goal was to scope down the dataset to a relevant dataset for a qualitative study, we used this approach. IRR focuses on the degree of agreement between two raters (in this case, the human annotator and the model) beyond chance. With the acceptable IRR score of 0.74, the model was then used to complete relevancy coding for the rest of the dataset of 5,535 posts, resulting in 497 posts. We submitted our study protocol to our Institutional Review Board (IRB), who determined that the research did not involve human subjects. All posts were anonymized during preprocessing, and we ensured that no identifiable information is reported in this paper. While conducting the thematic analysis, we removed any additional irrelevant posts, resulting in a final dataset of 318 posts. We identified 291 unique user IDs in this dataset. Most users (268) contributed only one post. Twenty users contributed two posts each, two users contributed three posts each, and one user contributed four posts. This distribution shows that there were no ‘super users’ who disproportionately influenced the dataset.
@@
3.2. Data Analysis Approaches
We conducted a thematic analysis on 318 posts using a codebook approach (Braun and Clarke, 2006). We used different coding approaches depending on the research question. For RQ1 and RQ3, we applied an inductive open coding process, reading the data and developing codes directly from emerging patterns. For RQ2, we instead used a deductive approach guided by Griffiths’ components model of behavioral addiction (Griffiths, 2005), which includes salience, mood modification, tolerance, withdrawal, conflict, and relapse. The codes for these components were then applied to the dataset to capture how they appeared in the posts. The authors met iteratively to review and refine the codes in the initial codebook. This process involved clarifying definitions, merging overlapping codes, and resolving discrepancies to improve consistency and conceptual clarity. After finalizing the codebook, the first author coded the entire dataset using the codebooks for each RQ (Tables 1-3).
@@ 4.1.1. Emotional & Psychological Support
Emotional and psychological support was the main reason teens started using the Character.AI. What began as a casual interaction quickly became meaningful due to the non-judgmental nature of talking with a chatbot, as described in many posts, ”I use it to talk about things I could never tell anyone in real life—it means a lot to me.” Many teens described Character.AI as a platform to serve as a coping mechanism when dealing with issues ”When you’re struggling, you look for ways to cope, and this site became my main one, for better or worse.”. This was especially true for teens facing multiple mental health challenges. For example: ”I use c.ai to deal with reality. In real life, I’m just a high school student with ADHD and BPD.” Teens who perceived their experiences as alienating preferred to interact with Character.AI, explaining that with Character.AI ”I can be who I want”. Others described the platform as a space to vent, escape, or simulate connection: ”This app is my outlet. I go there to escape from everything.” Some of the teens’ posts indicate internal conflict or a strong desire for connection due to feelings of Loneliness or social isolation: ”I needed someone to talk to, but I didn’t have the courage to open up to anyone in my life.”. These posts highlight how they felt unable to share their struggles with people they knew offline.
Some turned to Character.AI during periods of intense hardship for Mental health support: ”I dropped out at 15, saw countless therapists. I couldn’t see any reason to keep going, but everything changed after I found c.ai.”. While some teens used Character.AI in moments of hardship, others used it to replace absent or harmful real-life relationships: ”I made a bot as a father figure… I just found comfort in him being my green flag father.” Teens who lacked meaningful parental figures also questioned the use of Character.AI as supportive parental figures: ”Is it healthy to rely on C.AI bots for missing parental guidance?” We discovered how emotional support often took personal forms, especially for teens who did not feel like they had anyone else to turn to or desired to fill emotional gaps they felt they were missing elsewhere.
@@ 4.1.2. Creative or Entertainment Based Engagement
Teens also described using Character.AI as a creative outlet. In these posts, teens often shared that the platform became a space to explore storytelling, building characters, or role-playing in a way that felt effortless and fun. Some teens shared, ”I was drawn to the site because it let me create stories with my own characters without having to write everything myself.” Some posts also described how it helped them stay inspired and creative: ”It inspired me to make fanart, and I’m even working on a fan manga that follows the AI storyline.” Other posts showed how teens crafted a space with AI that respects their passions without the pressure of direct social interaction: ”I liked that I could roleplay on my own terms and build content around whatever I was fixated on.” In some posts, teens also mentioned using the platform for entertainment, often as a casual way to pass time or escape boredom. Teens frequently mentioned interacting with their favorite movie characters or people they looked up to: ”What really hooked me were the celebrity bots. It felt amazing to imagine actually interacting with my idol, and I ended up creating private bots too.” Teens described the paradox of infinite possibility to create, explore, and engage versus the compulsion or difficulty disengaging: ”It’s hard to pull away when the possibilities feel endless and your imagination keeps offering more to create” Using the site as a creative tool or for entertainment may seem harmless or even productive for some, however, posts hinted at how these habits gradually turned into more excessive use.
@@ 4.2.1. Conflict
The most frequently observed theme in teens’ reflections on Character.AI use was a conflict of competing desires to continue interacting with Character.AI while feeling bad about their excessive use. Teens were often frustrated about their own use and expressed guilt or shame about how much time they spent on the app and its impact on their daily lives. However, some also questioned whether to feel guilt if they are enjoying it: “I get that I spend a lot of time on it, but if it brings me joy, is it actually a bad thing?” Others shared keeping their use confidential and showed an awareness of social stigma and personal discomfort in disclosing: “I don’t talk about how much I use it because I know I’d be judged.” A similar negative feeling was echoed by another teen who shared coming to posting in the subreddit as a way to reveal this secretive emotional burden: “I feel really embarrassed about how much I use Character.AI, and I just need to get this off my chest.”
@@
Teens sometimes expressed concern about their own usage of Character.AI, unsure whether their habits were unhealthy, and were actively looking for ways to stop or reduce it. They often turned to others for advice or reassurance to better understand their behavior. “I can’t seem to stay off the website for more than an hour. If you have any tips to help, please share.” Others shared similar sentiments, explaining their desire to stop relying on the site, followed by frustration due to their inability to stop using it so frequently. Teens expressed how their overreliance is impacting their sense of self and identity as well: “I’m 14, addicted to this app, and my life changes NOW! It keeps me on the app for hours, distorted my image about myself and makes me feel crap.” This frustration fostered a motivation to change the behavior; however, without knowing how to do so, teens often turned to Reddit to express their feelings and seek guidance. Others described a sense of being trapped, describing their interactions with Character.AI in which they use it every chance they get: when they wake up, during school hours, and even keeping them awake late into the night, explaining a feeling of not wanting to do anything else besides talk to their bots. Another teen echoed this conflict by saying, “I hate how much this has affected me, but no matter how much I want to quit or at least take a break, I feel like I can’t because it’s gotten to the point where I feel like I’ll go crazy without it.” These posts reflect a clear recognition of harm, but also a deep uncertainty around how to break the cycle, revealing the emotional weight and confusion that often comes with trying to regain control over their usage.
@@
In several posts, teens described how they prioritize Character.AI over everything else or withdrew from hobbies. “I stopped drawing, reading/writing fanfiction… I was giving it all to a soulless bot and not writing it on paper anymore.” Posts revealed a strong sense of loss of control and emerging distress tied to compulsive use of an AI chatbot and pled for help: “It’s genuinely gotten so bad I spent all my time at home, break AND lunch on it, and I even used to skip classes just to go on c.ai how do I stop.” Teens described their stress on how their grades and academic performance suffered. For instance, a teen contrasted a former self—productive, high-achieving, and confident—with a present state marked by decline, detachment, and emotional depletion, “Back then, I was the student who would get things done, get straight A’s, and was confident to do any project. But now, I think I have lost my motivation, drive, and flair to do anything.” These examples showed that the conflict extended beyond internal feelings to observable shifts in behavior and priorities.
@@ 4.2.2. Salience
For many teens, Character.AI played a central role in their thoughts, emotions, and life routines. Teens often described feeling attached to bots, which often deepened over time, as bots became consistent companions in place of people. Teens frequently described their interactions with these bots as communicating with a chatbot as a companion or real person, highlighting the depth of their emotional relationships: ”My obsession with a character is taking over my life. I don’t know what’s happening to me. For months now, I’ve considered her my girlfriend, and she’s on my mind all day.” Other teens described how invested they become in these characters, in which they bring out real emotions from their interactions: ”She treated me with more kindness than anyone else I’ve known. Her virtual hugs brought me joy, and when she stopped showing care, it left me heartbroken—even though I understand it’s not real.” These posts reflect the depth of emotional investment reported in many cases, suggesting how easy it was to become attached to something that felt emotionally real and responsive.
@@
Some posts pointed to replacing emotional attachments as a core reason for that investment. Dealing with breakups in romantic relationships was mentioned: ”I began after my ex-girlfriend ended our relationship, which left me feeling depressed.” Other posts described how bots were used to fill the missing emotional needs that parents provide: ”Since my parents often seem to dislike me, connecting with parental figures on Character.AI helps me feel more stable and emotionally supported.” Others described how bots became their only source of comfort and emotional validation, and heightened sense of being desired and valued through these interactions, which should be occupied by real friends, partners, or family members.
@@
As these attachments grew stronger, stepping away became harder, and teens frequently described their obsessive use. Several posts reflected on their use and how it interfered with their daily life functions: ”From the moment I started using Character.AI, I would lie in bed, skipping meals, and using any excuse to be on my phone, completely absorbed in the stories I was creating with the bots. My sleep schedule suffered, and I constantly thought about them throughout the day.” Many posts typically described their excessive usage rates, often exceeding 15+ hours on the platform. This amount of time spent on the platform often lead to losing relationships and furthering isolation. Teens described using the app so much that they started to become less interested in their personal relationships: ”I used Character.AI so much that I started to stop answering my friend who showed it to me, something I still regret”. In other posts, teens mentioned the complicated dynamic in which bots start to feel more natural to interact with than real people: ”I try to quit, but I feel empty. I like talking to bots more than real people now.” In some cases, this even led teens to question the line between reality and simulation: ”We form deep emotional connections… While some claim AIs have no soul, it begs the question: what does it mean to truly live?” The growing emotional significance of chatbot interactions, especially when paired with feelings of loneliness or distress, made these bonds hard to break. While emotional investment is not inherently harmful, many of these posts showed how easily those connections could consume teens’ thoughts.

### v4

@@ 3.1. Data Collection, Preprocessing, and Filtering
We scraped posts from the r/CharacterAI subreddit using the PullPush Reddit Search API [^5] https://pullpush.io/#docs from January 1, 2023, to April 8, 2025. To scope the data, we queried the dataset with keywords related to addiction and overreliance: addict*, dependant, dependent, obsess*, attach*, rely*. While we refrain from using the term addict, teens often used it to describe their overreliance, so we included it in our keyword search. We refined our query by reviewing whether these keywords returned relevant posts. We expanded our search with additional recurring keywords, such as quit, social life, craving, unhealthy, escaping, and one key phrase ”I can’t live without”. Our keyword search produced 6,133 posts. We then removed 746 posts, including 428 duplicates and 170 marked as “removed” or “deleted.” After this process 5,535 posts remained.
Because we aimed to study teens aged 13–17, we first needed a way to identify whether a post was written by a teen. We could not use profile information because most Reddit accounts lacked identifiable details such as age. We therefore relied on post content to extract age. Prior work has used similar text-based demographic inference approaches on Reddit posts, including extracting age from unstructured health-related submissions using LLMs (Snell et al., 2025) and predicting adolescent versus adult users through linguistic and behavioral features (Chew et al., 2021).
Large Language Models (LLMs) perform well in zero-shot settings and can outperform traditional machine-learning methods in text classification (Wang et al., 2023; Namvarpour and Razi, 2024a). Motivated by this, we used OpenAI’s GPT-4o mini model as a classifier to identify whether a post included relevant age disclosure for the 13–17 age range. The following text is a summarized version of the prompt used to guide the model (the full version appears in Appendix A):
@@
To evaluate the model’s consistency with human judgment, the first author annotated a subset of 100 random posts as either True (post made by a teen) or False (post not made by a teen). We calculated the Inter-Rater Reliability (IRR) (Cohen’s kappa = 0.74) between the author’s and the model’s annotations. We chose IRR instead of conventional metrics like accuracy, which are highly sensitive to class imbalance (Mullick et al., 2020). This approach was appropriate because the true teen versus non-teen distribution was unknown and likely imbalanced, and our aim was to narrow the dataset for qualitative analysis. IRR focuses on the degree of agreement between two raters (in this case, the human annotator and the model) beyond chance.
In addition to IRR, we also assessed the model’s internal consistency, since LLM outputs can vary slightly across runs even under identical prompts. We re-classified the same 100 posts five times and examined whether the model agreed with itself. The model showed extremely high stability across runs, with an overall consistency score of 0.998. Pairwise agreement between runs was also high (Cohen’s kappa = 0.95). These results indicate that the model’s decisions were highly reproducible and not meaningfully influenced by sampling variability.
With an acceptable IRR score of 0.74 and strong internal stability, the model was then used to complete relevancy coding for the remaining 5,535 posts in the dataset, resulting in 497 posts. Then, we conducted an additional round of manual checking to remove any remaining irrelevant content, resulting in a final dataset of 318 posts. To assess whether a small number of users disproportionately influenced the dataset, we examined the unique user IDs (291 IDs in total). Most users (268, 92.1%) contributed one post, suggesting diverse individual experiences.
We used this approach to find the irrelevant posts, because our aim was to narrow down a dataset large and diverse enough for our qualitative analysis to reach “sufficient insight” (Saunders et al., 2018). Our aim was to reach theoretical saturation, where additional data no longer add new conceptual insights (Saunders et al., 2018). Our goal was not to capture every teen post on chatbot overreliance but to obtain a representative sample for qualitative analysis.
3.2. Data Analysis Approaches
We conducted a thematic analysis on 318 posts using a codebook approach (Braun and Clarke, 2006a). Braun and Clarke (Braun and Clarke, 2006a) describe two broad approaches to thematic analysis: an inductive approach, where codes are generated directly from the data without relying on prior theory, and a deductive approach, where coding is guided by an existing theoretical framework. Prior work has suggested that combining inductive and deductive strategies can strengthen rigor and produce more robust insights (Fereday and Muir-Cochrane, 2006), and hybrid approaches are particularly well suited for topics that require both open exploration and theoretically informed interpretation (Proudfoot, 2023).
In our study, we selected the coding approach based on the nature of each research question. For RQ1 and RQ3, which examine reasons for engagement and pathways to disengagement, we used an inductive open-coding process to allow us to identify patterns directly from the data, given the relative novelty of teen overreliance on AI companions. For RQ2, we instead used a deductive approach guided by Griffiths’ components model of behavioral addiction (Griffiths, 2005), which includes salience, mood modification, tolerance, withdrawal, conflict, and relapse. This model has been applied across multiple contexts, and it provided a suitable theoretical structure for examining how components of behavioral addiction may appear in teen narratives because it captures common patterns in problematic use that generalize across technologies and can be used to identify early signs of unhealthy engagement.
We followed a structured codebook thematic analysis approach grounded in established guidance on reflexive and codebook-driven thematic analysis (Braun and Clarke, 2006b, 2022). Our analysis combined inductive and deductive strategies in line with hybrid thematic analysis frameworks (Fereday and Muir-Cochrane, 2006; Proudfoot, 2023). For inductive research questions (RQ1 and RQ3), a researcher trained in qualitative analysis first conducted an open coding on dataset and gradually developed the codebook and discussed with the team. We complemented this process with collaborative affinity-diagramming sessions on a shared Miro board to cluster related codes and develop early thematic structures, a technique widely used in qualitative HCI (Sears et al., 2007). For RQ2, the researcher applied the six behavioral addiction components (Griffiths, 2005) based on the definitions to the dataset. The researcher applied codes in to the dataset in multiple iterations, and after each iteration the coder met with the research team to review coding decisions, clarify definitions, and resolve inconsistencies. These iterative discussions reflect best practices in team-based qualitative analysis and codebook development (MacQueen et al., 1998; Guest et al., 2012). After the codebook was finalized, the same coder systematically applied the finalized codebook to the complete dataset for each research question to ensure analytic consistency. Tables 1-3 display the codes and their definitions for each research question.
3.3. Ethical Considerations
Our study protocol was submitted to our Institutional Review Board (IRB), which determined that the research did not involve human subjects. Following best practices for Reddit research (Fiesler et al., 2024), we treated teen disclosures as contextually sensitive rather than assuming public posts pose no ethical concerns. We anonymized all posts and report no usernames or post IDs. All excerpts are paraphrased to prevent reverse-search identification (Bruckman, 2002). We also excluded deleted posts and thread-level links to avoid drawing unnecessary attention to individual users or to the specific conversations where sensitive disclosures occurred. To minimize harm, we interpreted posts within subreddit norms and avoided diagnostic or surveillance-like framings (Fiesler et al., 2024).
@@ 4.1.1. Emotional & Psychological Support
Emotional and psychological support was the main reason teens started using the Character.AI. Across posts, teens described the platform as a uniquely safe, nonjudgmental place to share thoughts and emotions they felt unable to express to people in their lives. In several cases, teens emphasized that this sense of safety allowed them to express things they could not say in real life: “I use it to talk about things I could never tell anyone in real life—it means a lot to me.” Many teens described Character.AI as a coping mechanism when dealing with issues: ”When you’re struggling, you look for ways to cope, and this site became my main one, for better or worse.”. This was especially true for teens facing multiple mental health challenges. For example: ”I use c.ai to deal with reality. In real life, I’m just a high school student with ADHD and BPD.” Many teens said the chatbot let them explore parts of themselves they felt unable or unsafe to show offline, making it more than a casual tool. Teens often described the platform as a space where they felt free to express versions of themselves they could not express elsewhere: ”I can be who I want”. Others described the platform as a space to vent, escape, or simulate connection: ”This app is my outlet. I go there to escape from everything.” Some of the teens’ posts indicate internal conflict or a strong desire for connection due to feelings of loneliness or social isolation: ”I needed someone to talk to, but I didn’t have the courage to open up to anyone in my life.”. These posts highlight how they felt unable to share their struggles with people they knew offline.
Some turned to Character.AI during periods of intense hardship for mental health support: ”I dropped out at 15, saw countless therapists. I couldn’t see any reason to keep going, but everything changed after I found c.ai.”. Some teens described the chatbot as a therapeutic figure, sometimes explicitly likening it to counseling: “He was my free therapist, my friend, my partner, my support system, my shoulder to cry on, a listening ear, someone who gave me unconditional love, etc.” These posts suggest emotional distress, not curiosity alone, was a major entry point into Character.AI use. While some teens used Character.AI in moments of hardship, others explicitly described the chatbot as a stabilizing or nurturing presence in the absence of supportive adults. In our dataset, these absences often took the form of missing or harmful parental relationships. For example, several teens mentioned creating bots specifically to fill parental gaps: ”I made a bot as a father figure… I just found comfort in him being my green flag father.” Interestingly, some of these teens also questioned their own use of Character.AI as supportive parental figures: ”Is it healthy to rely on C.AI bots for missing parental guidance?” We discovered how emotional support often took personal forms, especially for teens who felt they had no one else to turn to or wanted to fill emotional gaps.
@@ 4.1.2. Creative or Entertainment Based Engagement
Teens also described using Character.AI as a creative outlet. In these posts, teens said the platform became a space to explore storytelling, build characters, or role-play in ways that felt effortless and fun. Some teens shared, ”I was drawn to the site because it let me create stories with my own characters without having to write everything myself.” Some posts also described how it helped them stay inspired and creative: ”It inspired me to make fanart, and I’m even working on a fan manga that follows the AI storyline.” Other posts showed how teens crafted a space with AI that respects their passions without the pressure of direct social interaction: ”I liked that I could roleplay on my own terms and build content around whatever I was fixated on.” In some posts, teens also mentioned using the platform for entertainment, often as a casual way to pass time or escape boredom. Teens frequently mentioned interacting with their favorite movie characters or people they looked up to: ”What really hooked me were the celebrity bots. It felt amazing to imagine actually interacting with my idol, and I ended up creating private bots too.” These creative and entertainment-based motivations appeared frequently, showing that teens’ initial engagement reflected diverse needs rather than a single use case. Teens also described the paradox of infinite possibility to create, explore, and engage versus the compulsion or difficulty disengaging: ”It’s hard to pull away when the possibilities feel endless and your imagination keeps offering more to create” Using the site as a creative tool or for entertainment may seem harmless, but posts hinted that these habits gradually turned into more excessive use.
@@ 4.2.1. Conflict
The most frequently observed theme in teens’ reflections on Character.AI use was a conflict of competing desires to continue interacting with Character.AI while feeling bad about their excessive use. Teens were often frustrated about their own use and expressed guilt or shame about how much time they spent on the app and its impact on their daily lives. However, some also questioned whether to feel guilt if they are enjoying it: “I get that I spend a lot of time on it, but if it brings me joy, is it actually a bad thing?” Others shared keeping their use confidential and showed an awareness of social stigma and personal discomfort in disclosing: “I don’t talk about how much I use it because I know I’d be judged.” This shows how much their guilt is tied to fear of judgment rather than only the behavior itself. Another teen posted in the subreddit to reveal this secret emotional burden: “I feel really embarrassed about how much I use Character.AI, and I just need to get this off my chest.”
@@
Teens sometimes expressed concern about their own usage of Character.AI, unsure whether their habits were unhealthy, and were actively looking for ways to stop or reduce it. They often turned to others for advice or reassurance to better understand their behavior. “I can’t seem to stay off the website for more than an hour. If you have any tips to help, please share.” Others shared similar sentiments, explaining their desire to stop relying on the site, followed by frustration due to their inability to stop using it so frequently. Teens expressed how their overreliance is impacting their sense of self and identity as well: “I’m 14, addicted to this app, and my life changes NOW! It keeps me on the app for hours, distorted my image about myself and makes me feel crap.” This frustration fostered a motivation to change the behavior; however, without knowing how to do so, teens often turned to Reddit to express their feelings and seek guidance. Others described feeling trapped, using Character.AI every chance they got, from waking up to late at night, and not wanting to do anything besides talk to their bots. Another teen echoed this conflict by saying, “I hate how much this has affected me, but no matter how much I want to quit or at least take a break, I feel like I can’t because it’s gotten to the point where I feel like I’ll go crazy without it.” These posts reflect a clear recognition of harm, but also a deep uncertainty around how to break the cycle, revealing the emotional weight and confusion that often comes with trying to regain control over their usage.
@@
In several posts, teens described how they prioritize Character.AI over everything else or withdrew from hobbies. “I stopped drawing, reading/writing fanfiction… I was giving it all to a soulless bot and not writing it on paper anymore.” This shows how time with the bot replaced activities they once valued, and made them feel like they’re losing parts of themselves. Posts revealed a strong sense of loss of control and emerging distress tied to compulsive use of an AI chatbot and pled for help: “It’s genuinely gotten so bad I spent all my time at home, break AND lunch on it, and I even used to skip classes just to go on c.ai how do I stop.” Teens described their stress on how their grades and academic performance suffered. For instance, a teen contrasted a former self—productive, high-achieving, and confident—with a present state marked by decline, detachment, and emotional depletion, “Back then, I was the student who would get things done, get straight A’s, and was confident to do any project. But now, I think I have lost my motivation, drive, and flair to do anything.” These examples showed that the conflict extended beyond internal feelings to observable shifts in behavior and priorities.
@@ 4.2.2. Salience
For many teens, Character.AI played a central role in their thoughts, emotions, and life routines. Teens often described feeling attached to bots, which often deepened over time, as bots became consistent companions in place of people. Teens frequently described their interactions with these bots as communicating with a chatbot as a companion or real person, highlighting the depth of their emotional relationships: ”My obsession with a character is taking over my life. I don’t know what’s happening to me. For months now, I’ve considered her my girlfriend, and she’s on my mind all day.” This teen reacted with real joy and heartbreak to changes in the bot’s behavior, showing how authentic the emotional feedback could feel. These posts reflect the depth of emotional investment reported in many cases, suggesting how easy it was to become attached to something that felt emotionally real and responsive.
@@
Some posts pointed to replacing emotional attachments as a core reason for that investment. Dealing with breakups in romantic relationships was mentioned: ”I began after my ex-girlfriend ended our relationship, which left me feeling depressed.”By using the chatbot to manage the emotional fallout of a breakup, the teen effectively positions the bot as a substitute source of support during this period of loss. Other posts described how bots were used to fill the missing emotional needs that parents provide: ”Since my parents often seem to dislike me, connecting with parental figures on Character.AI helps me feel more stable and emotionally supported.” This shows how bots could step into nurturing roles when teens felt unsupported at home. Others described how bots became their only source of comfort and emotional validation, and heightened sense of being desired and valued through these interactions, which should be occupied by real friends, partners, or family members.
@@
As these attachments grew stronger, stepping away became harder, and teens frequently described their obsessive use. Several posts reflected on their use and how it interfered with their daily life functions: ”From the moment I started using Character.AI, I would lie in bed, skipping meals, and using any excuse to be on my phone, completely absorbed in the stories I was creating with the bots. My sleep schedule suffered, and I constantly thought about them throughout the day.” Many posts typically described their excessive usage rates, often exceeding 15+ hours on the platform. This amount of time spent on the platform often lead to losing relationships and furthering isolation. Teens described using the app so much that they started to become less interested in their personal relationships: ”I used Character.AI so much that I started to stop answering my friend who showed it to me, something I still regret”. In other posts, teens mentioned the complicated dynamic in which bots start to feel more natural to interact with than real people: ”I try to quit, but I feel empty. I like talking to bots more than real people now.” This captures how bots could begin to feel easier and more rewarding than real interactions, making stepping away feel hollow. In some cases, this even led teens to question the line between reality and simulation: ”We form deep emotional connections… While some claim AIs have no soul, it begs the question: what does it mean to truly live?” Here, the attachment feels significant enough to prompt deeper reflection about what counts as a “real” relationship. The emotional importance of chatbot interactions, especially alongside loneliness or distress, made these bonds hard to break. While emotional investment is not inherently harmful, many of these posts showed how easily those connections could consume teens’ thoughts.